Chlorine can be prepared in the laboratory by the reaction of manganese dioxide with hydrochloric acid, HCl(aq) , as described by the chemical equation MnO2(s)+4HCl(aq)⟶MnCl2(aq)+2H2O(l)+Cl2(g) How much MnO2(s) should be added to excess HCl(aq) to obtain 175 mL Cl2(g) at 25 °C and 785 Torr ?
Answer #1

PV = nRT

where, P = pressure = 785 torr = 1.03 atm

V = volume = 175 mL = 0.175 L

n = number of moles of Cl2

R = Gas constant

T = temperature = 25 + 273 = 298 K

1.03 * 0.175 = n * 0.0821 * 298

0.180 = n * 24.5

n = 0.180 / 24.5 = 0.00735 mol

Therefore, the number of moles of Cl2 = 0.00735 mol

From the balanced equation we can say that

1 mole of Cl2 produced by 1 mole of MnO2 so

0.00735 mole of Cl2 will be produced by

= 0.00735 mole of Cl2 *(1 mole of MnO2 / 1 mole of Cl2)

= 0.00735 mole of MnO2

mass of 1 mole of MnO2 = 86.9368 g

so the mass of 0.00735 mole of MnO2 = 0.639 g

Therefore, the mass of MnO2 required would be 0.639 g